# Galled Perfect Transfer Networks

### v2

@@ 4 The Galled Compatibility Problem
Let us recall the galled compatibility problem: we are given a set of characters $\mathcal{C}$ on a set of taxa $\mathcal{S}$ and must decide whether a galled tree $N$ explains $\mathcal{C}$. Note that if such an $N$ exists, then its base tree is galled-completable. Therefore, $\mathcal{C}$ is galled-compatible if and only if there exists a tree $T$ that is galled-completable for $\mathcal{C}$. Instead of aiming to construct $N$ directly, our strategy is to build such a $T$ using the characterizations from the previous section.
@@
Figure 5: An example instance with characters $\mathcal{C}=\{A,B,C,X,Y,P,Q,R\}$ on taxa $\mathcal{S}=\{a,b,c,d,e,f,g,h,i,j,k,l,m\}$ used to illustrate the main steps 1-4 of the algorithm. Step (1) splits $A$ into $A\setminus B,A\cap B$. Step (2) integrates the clades forced by $X,Y$, which intersect the two clades of $A$. Step (3) integrates $B$ and $C$ as clades. Step (4) solves for $P$ recursively, and then for $Q,R$ recursively, then the transfer required by $A,X$, and $Y$ is added.
The detailed algorithm is somewhat involved, but the main idea can be described in a few steps that are illustrated in Figure 5:
@@
(1) otherwise, every maximal character $A$ has some incompatibility with some $B$, as in Figure 5. We can show that $T$ must either split $A$ into clades $\{A\setminus B,A\cap B\}$, or split $B$ into clades $\{B\setminus A,B\cap A\}$. We try the first option, and if it leads to a dead-end we try the other option. That is, for each of these two possibilities, we initiate a tree with only the two clades and proceed to the next steps. For the rest of the description, we assume that we start a tree with $A\setminus B,A\cap B$ as in Figure 5.1. This means that in any completion, there will be a transfer edge between the two clades to provide an origin for $A$;
(2) each character that intersects both clades must use that same transfer edge in a completion. Such characters enforce new clades that refine the previous ones. For example in Figure 5.2, $X$ enforces the clade $\{d,e\}$, which is the intersection $X\cap\{a,b,c,d,e\}$, and similarly $Y$ enforces $\{c,d,e\}$. These enforced new clades must be ordered by inclusion;
(3) other characters can enforce further clades, namely those that contain the minimal clades enforced so far. For example, the character $C$ from Figure 5 contains $\{d,e\}$, and so the clade $C$ is enforced. Likewise, $B$ is forced since it contains $\{f,g,h\}$. These are added in Figure 5.3;
(4) it turns out that if $\mathcal{C}$ is galled-compatible, then any unforced character so far represents a set of leaves that have the same parent $v$ in $T$. See $P,Q,R$ in Figure 5.3. We can recurse into the leaf children of each $v$ and replace the leaves by the resulting network. In Figure 5.4, the leaf set $\{i,j,k\}$ is replaced by a network that explains $Q,R$, and $\{f,g,h\}$ by a network that explains $P$.
@@
If $C$ as stated exists, we can start constructing a tree whose root has two children, one with $C$ as a clade, and the other with everything not in $C$. This automatically explains $C$. We can then recursively solve for $\mathcal{C}_{1}$ on the $C$ side, and for $\mathcal{C}_{2}$ on the non-$C$ side, and replace the child clades with the corresponding networks. Because $C$ is compatible, every character will be on one side, this solves for every character. ∎
@@
By incompatibility, $A$ and $B$ cannot both be clades of $T$. If only one of them is a clade, say $A$, then we are done since the only way to have two FAs for $B$ is to put $B\cap A$ inside of $A$, and $B\setminus A$ elsewhere. So assume that both $A$ and $B$ are split, having two FAs $a_{1},a_{2}$ and $b_{1},b_{2}$ each. Because $A$ and $B$ intersect, with some effort it can be shown that these FAs must be related by ancestry (i.e., $a_{1}\prec b_{1}$ or $b_{1}\prec a_{1}$, and also $a_{2}\prec b_{2}$ or $b_{2}\prec a_{2}$). The proof shows that each possible case leads to requiring two distinct transfers to explain $A$ and $B$, which create intersecting cycles. ∎
@@
We claim that some $a_{i}$ is a strict ancestor of some $b_{j}$ or vice-versa. Note that $A\cap B\neq\emptyset$, so we may assume without loss of generality that $A_{1}\cap B_{1}\neq\emptyset$. Thus, $a_{1}$ has a descending leaf in common with $b_{1}$, which implies that $a_{1}$ is an ancestor of $b_{1}$ or vice-versa. If $a_{1}\neq b_{1}$, our claim holds. If $a_{1}=b_{1}$, then $a_{2}\prec b_{2}$ or $b_{2}\prec a_{2}$ by Lemma 3 (part 2), and again our claim holds.
We may therefore assume, without loss of generality, that $a_{1}$ is a strict ancestor of $b_{1}$, which implies $B_{1}\subseteq A_{1}$. Notice that $b_{2}$ cannot be a descendant of $a_{1}$, since otherwise we would have $B_{2}\subseteq A_{1}$ and $B_{1}\cup B_{2}\subseteq A_{1}$, contradicting the incompatibility of $A$ and $B$. Since $b_{2}$ is incomparable with $b_{1}$, it also cannot be an ancestor of $a_{1}$, and thus $b_{1}$ is also incomparable with $a_{1}$. Then by Lemma 3 (part 1), $a_{2}=b_{2}$. Those imply $B_{1}\subseteq A_{1}$ and $A_{2}=B_{2}$, in turn implying $B\subseteq A$ and contradicting that $A$ and $B$ are incompatible.
This establishes that exactly one of $A$ or $B$ is a clade of $T$. Suppose that $B$ is a clade and that $A$ is split into $A_{1}$ and $A_{2}$ in $T$, with respective FAs $a_{1},a_{2}$. Let $b$ be the (unique) FA of $B$. If $b\preceq a_{1}$ or $b\preceq a_{2}$, then $B\subseteq A$ and $A,B$ would be compatible. Moreover, $B$ intersects with at least one of $A_{1}$ and $A_{2}$ since it intersects with $A$, and thus $b$ is an ancestor of $a_{1}$ or $a_{2}$. If $b$ is an ancestor of both, then $A\subseteq B$ and again $A,B$ would be compatible. Hence $b$ is an ancestor of exactly one of $a_{1}$ or $a_{2}$. If $b$ is an ancestor of $a_{1}$, then $A_{1}=B\cap A$ and $A_{2}=A\setminus B$. If $b$ is an ancestor of $a_{2}$, then $A_{2}=B\cap A$ and $A_{1}=A\setminus B$. This shows that the first case of the statement holds.
@@
We call $X_{1}\cap A_{1}$ the bottom of the chain, and we call $A_{2}$ the stable side of the chain. In Figure 5.3, $A_{1}=\{a,b,c,d,e\},A_{2}=\{f,g,h\}$ and $\mathcal{X}=\{X,Y,A\}$. One can see that $\mathcal{X}$ forms an $(A_{1},A_{2})$-chain with bottom $\{d,e\}$ and stable side $\{f,g,h\}$. The concept of a chain is illustrated in Figure 7.
Figure 7: An $(A_{1},A_{2})$-chain, where $X_{1},X_{2},X_{3}\in\mathcal{X}$ all intersect with both $A_{1}$ and $A_{2}$. On the $A_{1}$ side, the FAs are ordered by ancestry, and the only way to explain all these characters is to have them use the same transfer to reach the $A_{2}$ side.
@@
Then, after possibly exchanging the subscripts of $A_{1}$ and $A_{2}$, $\mathcal{X}$ is an $(A_{1},A_{2})$-chain. Moreover, for every $X\in\mathcal{X}$, the clades $X\cap A_{1}$ and $X\cap A_{2}$ are in $T$.
@@
Let $A$ be a maximal character of $\mathcal{C}$. Suppose that $T$ is a galled-completable tree for $\mathcal{C}$ in which $A$ is split into the clades $A_{1}$ and $A_{2}$. Let $\mathcal{X}\subseteq\mathcal{C}$ be the subset of characters that intersect with both $A_{1}$ and $A_{2}$ and suppose that $\mathcal{X}$ is an $(A_{1},A_{2})$-chain. Let $X_{1}\cap A_{1}$ be the bottom of the chain and let $A_{2}$ be the stable side of the chain.
@@
To put this precisely, let us gather all the information on the tree that has been shown to be forced so far. Let $A$ be a maximal character of $\mathcal{C}$ and let $\{A_{1},A_{2}\}$ be a partition of $A$ into two non-empty sets. Let $\mathcal{X}$ be the characters that intersect with $A_{1}$ and $A_{2}$ and suppose that $\mathcal{X}$ forms and $(A_{1},A_{2})$-chain. We say that a character $C\in\mathcal{C}$ is forced by $\{A_{1},A_{2}\}$ if either: $C\in\mathcal{X}$; $C$ contains the bottom of the $\mathcal{X}$ chain; or $C$ contains the stable side of the $\mathcal{X}$ chain.
Furthermore, a clade $Y\subseteq\mathcal{S}$ is forced by $\{A_{1},A_{2}\}$ if either: $Y=X\cap A_{1}$ or $Y=X\cap A_{2}$ for some $X\in\mathcal{X}$; or $Y=C$ for some character $C\in\mathcal{C}$ that contains the bottom or the stable side of the $\mathcal{X}$ chain. The first type of clades are enforced due to Lemma 4, and the second type due to Lemma 4.
@@
Figure 8: An illustration of Lemma 4. Left: the tree $T$ that contains the forced clades, with $X_{1}\cap A_{1}$ the bottom of the chain and $A_{2}$ the stable side. We assume that some characters $B,D,E$ enforced other clades. If we assume that some character $C$ has elements with distinct parents, the situation on the right is unavoidable. A transfer will be needed to link the bottom and stable side, and another transfer to explain $C$. These two transfers create intersecting underlying cycles.
@@
In the ()($\Rightarrow$)( ⇒ ) direction, we know by all the previous lemmata that all the stated conditions must be satisfied by a galled-completable tree for $\mathcal{C}$. In the ()($\Leftarrow$)( ⇐ ) direction, suppose that all the conditions hold. Let $T$ be a tree that contains all the clades forced by $\{A_{1},A_{2}\}$. Then we can explain all the forced characters by adding a transfer edge just above the bottom of the $\mathcal{X}$ chain, going just above the stable side. As for the other characters, for a node $v$ of $T$, let $\mathcal{C}_{v}\subseteq\mathcal{C}$ be the non-forced characters having taxa whose parent is $v$. We can find a galled network $N_{v}$ that explains $\mathcal{C}_{v}$ and “attach” $N_{v}$ as a child of $v$. Doing this for every $v$ does not create intersecting cycles and explains all the remaining characters. ∎
@@
Suppose that $\mathcal{C}$ has no maximal character that is compatible with all the others. Then there exists a pair of incompatible characters $\{A,B\}$ of $\mathcal{C}$ such that $A$ and $B$ are both maximal.
@@
1 fonction getGalledTree($\mathcal{C}$)

### v3

@@ 4 The Galled Compatibility Problem
Let us recall the galled compatibility problem: we are given a set of characters $\mathcal{C}$ on a set of taxa $\mathcal{S}$ and must decide whether a galled LGT network $N$ explains $\mathcal{C}$. Note that if such an $N$ exists, then its base tree is galled-completable. Therefore, $\mathcal{C}$ is galled-compatible if and only if there exists a tree $T$ that is galled-completable for $\mathcal{C}$. Instead of aiming to construct $N$ directly, our strategy is to build such a $T$ using the characterizations from the previous section.
@@
Figure 5: An example instance with characters $\mathcal{C}=\{A,B,C,X,Y,P,Q,R\}$ on taxa $\mathcal{S}=\{a,b,c,d,e,f,g,h,i,j,k,l,m\}$ used to illustrate the main steps 1-4 of the algorithm. Note that the maximal characters are: $A,B,C,Q$ and $R$. Step (1) splits $A$ into $A\setminus B,A\cap B$. Step (2) integrates the characters $X,Y$ that intersect the two clades of $A$. Step (3) integrates $B$ and $C$ as clades. Step (4) solves for $P$ recursively, and then for $Q,R$ recursively. Note that transfer edges are not part of the output of the algorithm, as it only returns a galled-completable tree. The transfer edges were added in the subfigure to show that the resulting tree was indeed galled-completable — for instance the transfer edge from $\{d,e\}$ to $\{f,g,h\}$ is required to explain $A,X,Y$.
We now present our polynomial-time algorithm to reconstruct a galled-completable tree $T$ for $\mathcal{C}$, if one exists. For each character $C\in\mathcal{C}$, we must decide whether $C$ should be a clade in $T$, or whether we want to split $C$ into two different clades that will eventually be joined by a transfer edge in a completion of $T$. The detailed algorithm is somewhat involved, but the main idea can be described in a few steps that are illustrated in Figure 5:
@@
(1) otherwise, every maximal character $A$ has some incompatibility with some $B$, as in Figure 5. We choose an arbitrary pair of incompatible characters $\{A,B\}$ such that both $A$ and $B$ are maximal (such a pair is shown to exist).
We can show that $T$ must either split $A$ into clades $\{A\setminus B,A\cap B\}$ and keep $B$ as a clade, or split $B$ into clades $\{B\setminus A,B\cap A\}$ and keep $A$ as a clade. We try the first option, and if it leads to a dead-end we try the other option. That is, for each of these two possibilities, we initiate a tree with only the two clades and proceed to the next steps. For the rest of the description, we assume that we start a tree with $A\setminus B,A\cap B$ as in Figure 5.1. This means that in any completion, there will be a transfer edge between the two clades to provide an origin for $A$;
(2) We next scan for clades that are forced by the above choice, i.e., clades that must be present in any galled-completable tree that contains clades $A\setminus B,A\cap B$. We show that if a character intersects both these clades, then both intersections are forced clades. For example in Figure 5.2, $X$ enforces the clades $X\cap(A\setminus B)=\{d,e\}$ and $X\cap(A\cap B)=\{f,g,h\}$. Similarly $Y$ enforces $Y\cap(A\setminus B)=\{c,d,e\}$ and $Y\cap(A\cap B)=\{f,g,h\}$;
(3) We can find further forced clades. Namely, the characters that contain clades enforced so far are shown to be forced clades. For example, the character $C$ from Figure 5 is a superset of the clade $\{d,e\}$ enforced in the previous step, and so the clade $C$ is enforced. Likewise, $B$ is forced since it contains $\{f,g,h\}$. These are added in Figure 5.3;
(4) it turns out that if $\mathcal{C}$ is galled-compatible, then any character that has not implied a forced clade so far represents a set of leaves that have the same parent $v$ in $T$. See $P,Q,R$ in Figure 5.3. We can recurse into the leaf children of each $v$ and replace the leaves by a galled-completable subtree with respect to that set of leaves. In Figure 5.4, the leaf set $\{i,j,k\}$ is replaced by a tree that is galled-completable for $Q,R$, and $\{f,g,h\}$ by a tree that is galled-completable for $P$.
@@
The forward direction is immediate. For the converse, if $C$ as stated exists, we can start constructing a tree whose root has two children, one with $C$ as a clade, and the other with everything not in $C$. This automatically explains $C$. We can then recursively solve for $\mathcal{C}_{1}$ on the $C$ side, and for $\mathcal{C}_{2}$ on the non-$C$ side, and replace the child clades with the corresponding networks. Because $C$ is compatible with every character, they will all be in either $\mathcal{C}_{1}$ or $\mathcal{C}_{2}$. If the latter two are galled-compatible, we can easily merge the networks that explains each subset as in Figure 6, since characters in $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ have no taxa in common. ∎
@@
By incompatibility, $A$ and $B$ cannot both be clades of $T$. If only one of them is a clade, say $A$, then we are done since the only way to have two FAs for $B$ is to put $B\cap A$ inside of $A$, and $B\setminus A$ elsewhere. So assume that both $A$ and $B$ are split, having two FAs $a_{1},a_{2}$ and $b_{1},b_{2}$ each. Because $A$ and $B$ intersect, with some effort it can be shown that these FAs must be related by ancestry (i.e., $a_{1}\prec_{T}b_{1}$ or $b_{1}\prec_{T}a_{1}$, and also $a_{2}\prec_{T}b_{2}$ or $b_{2}\prec_{T}a_{2}$). The proof shows that each possible case leads to requiring two distinct transfers to explain $A$ and $B$, which create intersecting cycles. ∎
@@
We claim that in $T$, some $a_{i}$ is a strict ancestor of some $b_{j}$ or vice-versa. Note that $A\cap B\neq\emptyset$, so we may assume without loss of generality that $A_{1}\cap B_{1}\neq\emptyset$. Thus, $a_{1}$ has a descending leaf in common with $b_{1}$, which implies that $a_{1}$ is an ancestor of $b_{1}$ or vice-versa. If $a_{1}\neq b_{1}$, our claim holds. If $a_{1}=b_{1}$, then $a_{2}\prec_{T}b_{2}$ or $b_{2}\prec_{T}a_{2}$ by Lemma 3 (part 2), and again our claim holds.
We may therefore assume, without loss of generality, that $a_{1}$ is a strict ancestor of $b_{1}$, which implies $B_{1}\subseteq A_{1}$. Notice that $b_{2}$ cannot be a descendant of $a_{1}$, since otherwise we would have $B_{2}\subseteq A_{1}$ and $B_{1}\cup B_{2}\subseteq A_{1}$, contradicting the incompatibility of $A$ and $B$. Since $b_{2}$ is incomparable with $b_{1}$, it also cannot be an ancestor of $a_{1}$, and thus $b_{2}$ is also incomparable with $a_{1}$. Then by Lemma 3 (part 1), $a_{2}=b_{2}$. Those imply $B_{1}\subseteq A_{1}$ and $A_{2}=B_{2}$, in turn implying $B\subseteq A$ and contradicting that $A$ and $B$ are incompatible.
This establishes that exactly one of $A$ or $B$ is a clade of $T$. Suppose that $B$ is a clade and that $A$ is split into $A_{1}$ and $A_{2}$ in $T$, with respective FAs $a_{1},a_{2}$. Let $b$ be the (unique) FA of $B$. If $b\preceq_{T}a_{1}$ or $b\preceq_{T}a_{2}$, then $B\subseteq A$ and $A,B$ would be compatible. Moreover, $B$ intersects with at least one of $A_{1}$ and $A_{2}$ since it intersects with $A$, and thus $b$ is an ancestor of $a_{1}$ or $a_{2}$. If $b$ is an ancestor of both, then $A\subseteq B$ and again $A,B$ would be compatible. Hence $b$ is an ancestor of exactly one of $a_{1}$ or $a_{2}$. If $b$ is an ancestor of $a_{1}$, then $A_{1}=B\cap A$ and $A_{2}=A\setminus B$. If $b$ is an ancestor of $a_{2}$, then $A_{2}=B\cap A$ and $A_{1}=A\setminus B$. This shows that the first case of the statement holds.
@@
We call $X_{1}\cap A_{1}$ the bottom of the chain, and we call $A_{2}$ the stable side of the chain.
Consider Figure 5.3, with character $A$ partitioned into $A_{1}=\{a,b,c,d,e\},A_{2}=\{f,g,h\}$ and $\mathcal{X}=\{X,Y,A\}$. One can see that $\mathcal{X}$ forms an $(A_{1},A_{2})$-chain. Indeed, we have $X\cap A_{1}\subset Y\cap A_{1}\subset A\cap A_{1}$ because these intersections give $\{d,e\}\subset\{c,d,e\}\subset\{a,b,c,d,e\}$. Moreover, $X\setminus A_{1}=Y\setminus A_{1}=A\setminus A_{1}=\{f,g,h\}$. The bottom of this chain is $\{d,e\}$ and the stable side $\{f,g,h\}$. The general concept of a chain is illustrated in Figure 7 (the transfer edge can be ignored).
Figure 7: An $(A_{1},A_{2})$-chain, where $X_{1},X_{2},X_{3}\in\mathcal{X}$ all intersect with both $A_{1}$ and $A_{2}$. On the $A_{1}$ side, the FAs are ordered by ancestry, and the only way to explain all these characters is to have them use the same transfer to reach the $A_{2}$ side. Note, the transfer edge is not part of the chain nor the reconstruction, we add it to emphasize how the characters forming the chain can be explained.
@@
Then, after possibly exchanging the subscripts of $A_{1}$ and $A_{2}$, $\mathcal{X}$ is an $(A_{1},A_{2})$-chain. Moreover, for every $X\in\mathcal{X}$, the clades $X\cap A_{1}$ and $X\cap A_{2}=A_{2}$ are in $T$.
@@
Let $A$ be a maximal character of $\mathcal{C}$. Suppose that $T$ is a galled-completable tree for $\mathcal{C}$ in which $A$ is split into the clades $A_{1}$ and $A_{2}$. Let $\mathcal{X}\subseteq\mathcal{C}$ be the subset of characters that intersect with both $A_{1}$ and $A_{2}$ and suppose that $\mathcal{X}$ is a $(A_{1},A_{2})$-chain. Let $X_{1}\cap A_{1}$ be the bottom of the chain and let $A_{2}$ be the stable side of the chain.
@@
To put this precisely, let us gather all the information on the tree that has been shown to be forced so far.
Definition 5 (Forced characters and clades.).
Let $A$ be a maximal character of $\mathcal{C}$ and let $\{A_{1},A_{2}\}$ be a partition of $A$ into two non-empty sets. Let $\mathcal{X}$ be the characters that intersect with $A_{1}$ and $A_{2}$ and suppose that $\mathcal{X}$ forms and $(A_{1},A_{2})$-chain. We say that a character $C\in\mathcal{C}$ is forced by $\{A_{1},A_{2}\}$ if either: $C\in\mathcal{X}$; $C$ contains the bottom of the $\mathcal{X}$ chain; or $C$ contains the stable side of the $\mathcal{X}$ chain.
Furthermore, a clade $Y\subseteq\mathcal{S}$ is forced by $\{A_{1},A_{2}\}$ if either: $Y=X\cap A_{1}$ or $Y=X\cap A_{2}$ for some $X\in\mathcal{X}$; or $Y=C$ for some character $C\in\mathcal{C}$ that contains the bottom or the stable side of the $\mathcal{X}$ chain.
For example in Figure 5.2, we made the decision of partitioning $A$ into $A_{1}$ and $A_{2}$, and $\mathcal{X}=\{X,Y,A\}$ forms an $(A_{1},A_{2})$-chain. The definition states that all characters of $\mathcal{X}$ are forced. Moreover, character $C$ is forced because it is a superset of the bottom of the chain $\{d,e\}$, and character $B$ is forced because it contains the stable side $\{f,g,h\}$. We call these characters forced because by Lemma 4 and Lemma 4, they imply the existence of clades in $T$. In turn, the clades that are implied are called forced. For example, character $X$ in the figure implies the existence of the forced clade $X\cap A_{1}=\{d,e\}$ in $T$, and character $C$ implies the existence of forced clade $C$ in $T$.
@@
Figure 8: An illustration of Lemma 4. Left: the tree $T$ that contains the forced clades, with $X_{1}\cap A_{1}$ the bottom of the chain and $A_{2}$ the stable side. We assume that some characters $B,D,E$ enforced other clades. If we assume that some character $C$ has two taxa with distinct parents (gold circles in the figure), the situation on the right is unavoidable. That is, in any galled-completion of $T$, a transfer will be needed to link the bottom and stable side, and another transfer to explain $C$. These two transfers create intersecting underlying cycles.
@@
In the ()($\Rightarrow$)( ⇒ ) direction, we know by all the previous lemmata that all the stated conditions must be satisfied by a galled-completable tree for $\mathcal{C}$. In the ()($\Leftarrow$)( ⇐ ) direction, suppose that all the conditions hold. Let $T$ be a tree that contains all the clades forced by $\{A_{1},A_{2}\}$. Then we can explain all the forced characters by adding a transfer edge just above the bottom of the $\mathcal{X}$ chain, going just above the stable side. As for the other characters, for a node $v$ of $T$, let $\mathcal{C}_{v}\subseteq\mathcal{C}$ be the non-forced characters having taxa whose parent is $v$. We can find a galled LGT network $N_{v}$ that explains $\mathcal{C}_{v}$ and “attach” $N_{v}$ as a child of $v$. Doing this for every $v$ does not create intersecting cycles and explains all the remaining characters. ∎
@@
Suppose that $\mathcal{C}$ has no maximal character that is compatible with all the other characters. Then there exists a pair of incompatible characters $\{A,B\}$ of $\mathcal{C}$ such that $A$ and $B$ are both maximal.
@@
1 function getGalledTree($\mathcal{C}$)